OpenStudy (anonymous):

order of operations to simplify 13/6 - 5/18 X 3

OpenStudy (anonymous):

13/6 - 5/6 = 8/6 =4/3

OpenStudy (anonymous):

did you do the multiply 3 in there?

OpenStudy (anonymous):

5/18*3 = 5/6

OpenStudy (anonymous):

so what is the final answer for 13/6 - 5/18 X(times) 3

OpenStudy (anonymous):

4/3